# SAE-V: Interpreting Multimodal Models for Enhanced Alignment

### v1

@@ Abstract
With the integration of image modality, the semantic space of multimodal large language models (MLLMs) is more complex than text-only models, making their interpretability more challenging and their alignment less stable, particularly susceptible to low-quality data, which can lead to inconsistencies between modalities, hallucinations, and biased outputs. As a result, developing interpretability methods for MLLMs is crucial for improving alignment quality and efficiency. In text-only LLMs, Sparse Autoencoders (SAEs) have gained attention for their ability to interpret latent representations. However, extending SAEs to multimodal settings presents new challenges due to modality fusion and the difficulty of isolating cross-modal representations. To address these challenges, we introduce SAE-V, a mechanistic interpretability framework that extends the SAE paradigm to MLLMs. By identifying and analyzing interpretable features along with their corresponding data, SAE-V enables fine-grained interpretation of both model behavior and data quality, facilitating a deeper understanding of cross-modal interactions and alignment dynamics. Moreover, by utilizing cross-modal feature weighting, SAE-V provides an intrinsic data filtering mechanism to enhance model alignment without requiring additional models. Specifically, when applied to the alignment process of MLLMs, SAE-V-based data filtering methods could achieve more than 110% performance with less than 50% data. Our results highlight SAE-V’s ability to enhance interpretability and alignment in MLLMs, providing insights into their internal mechanisms.
@@ 1 Introduction
With the development and success of large language models (LLMs) (Dubey et al., 2024; Achiam et al., 2023), researchers have begun to introduce visual understanding to these models, thereby extending their operational scope from language to a mix of vision and language, resulting in the creation of powerful multimodal large language models (MLLMs) (Alayrac et al., 2022; Liu et al., 2024; Team et al., 2024; Team, 2024). To enhance the multimodal understanding capabilities of MLLMs, the research community has explored various architectures, including using individual image/text encoders to encode cross-modal information into a joint representation space (Zhang et al., 2023; Liu et al., 2024; Zhu et al., 2024; Wu et al., 2024b) and leveraging image tokenizers to transform all inputs into a unified token sequence (Team, 2024; Xie et al., 2024; Wu et al., 2024a; Wang et al., 2024). Despite the difference in the architectures of these models, their essential goal is the same: Fuse the text and image representation space into a joint multimodal semantic space.
As MLLMs continue to scale up in both size and capability, their interpretability and controllability remain a significant challenge (Zhang & Zhu, 2018; Stan et al., 2024). Currently, mechanistic interpretability techniques such as circuit analysis (Olsson et al., 2022) and dictionary learning with sparse autoencoders (Cunningham et al., 2023) are the most widely recognized approaches to interpreting LLMs. However, their application to MLLMs, especially in the context of cross-modal integration, has been limited. There is a pressing need for specialized tools and frameworks that can unravel the intricate workings of MLLMs.
Moreover, current interpretability efforts are focused mainly on interpreting models, rather than applying this interpretability to real alignment situations, which also makes it difficult to evaluate these methods effectively. Top-down approaches, such as Representation Engineering (Zou et al., 2023) and activation steering (Turner et al., 2023; Panickssery et al., 2023), can directly evaluate the control effects of interpretability methods through control or unlearning techniques. However, for bottom-up methods like circuit analysis (Wang et al., 2022), sparse autoencoders (SAEs) (Cunningham et al., 2023), and cross-coders, effective evaluation methods beyond loss reduction are limited. Based on the previous discussion, can we propose a bottom-up multimodal interpretability approach that can directly enhance the alignment process?
@@ 2.1 Preliminary: Sparse Autoencoder Paradigm
We adopt SAE-V (denoted as $\mathcal{S}_{\theta}$) architecture from the methodology proposed in (Bricken et al., 2023), which comprises an encoder and a feature dictionary $\mathcal{F}_{\theta}\text{:}\ \{\bm{f}_{k}\}_{k=1}^{\bm{n}}$ as a decoder. Let the input be denoted as $H\in\mathbb{R}^{\bm{l}\times\bm{m}}$, the hidden state of a specific layer of a MLLM $\mathcal{M}_{\theta}$. The SAE-V encoding operation $\mathcal{S}_{\theta}(\cdot)$ is
$Z=\text{ReLU}(W_{\text{enc}}H+b_{enc}),$ (1)
@@
$\mathcal{L}_{R}={||H-Z\times(\bm{f}_{1},\bm{f}_{2},\dots,\bm{f}_{m})^{\top}||}_{2}^{2}.$ (2)
@@
[…]
@@
        $\mathcal{A}_{k}\leftarrow\mathcal{A}_{k}\cup\{\bm{h}_{j}\text{:}\ \bm{h}_{j}\in H_{i},\bm{z}_{j}=\bm{e}_{j}Z_{i},\bm{z}_{j}\bm{f}_{k}^{\top}>\delta\}$
@@
     $F_{i}\leftarrow\{\bm{f}_{k}\text{:}\|\bm{f}_{k}Z_{i}^{\top}\|_{\infty}>\delta\}$
@@ 2.2 Interpreting Multimodal Data with SAE-V
It has been previously demonstrated (Gao et al., 2024; Cunningham et al., 2023) that SAE can be employed to interpret how LLMs encode semantic information from these models. This feature motivates us to apply SAE-V to assess the quality of the data and thus facilitate data filter for alignment.
We adopt a cosine similarity score ranking algorithm for data filtering (shown in Algorithm 1). Let the multimodal training dataset be donated as $\mathcal{D}=\{\bm{d}_{i}\}_{i=1}^{p}$, where $\bm{d}_{i}=\{\bm{u}_{j}\text{:}\bm{u}_{j}\in\mathcal{T}\lor\bm{u}_{j}\in\mathcal{V}\}_{j=1}^{m}$ is a set of tokens from text vocabulary $\mathcal{T}$ and tokens from vision vocabulary $\mathcal{V}$. We acquire feature activation token $\bm{z}_{j}$ by MLLM forward and Equation 1, i.e.
@@
We define a SAE-V feature $\bm{f}_{k}$ is activated on $\bm{z}_{j}\ \textit{if}\ \bm{z}_{j}\bm{f}_{k}^{\top}>\delta$, where $\delta$ is activation bound. Correspondingly, we state that $\bm{f}_{k}$ is activated on $\bm{d}_{i}\ \textit{if}\ \exists\ \bm{z}_{j}\in Z_{i}\ \text{activating}\ \bm{f}_{k}$.
@@
    SAE-V Feature Weighting For each feature $\bm{f}_{k}$, we identify its top-K hidden state text tokens $\bm{t}=\textit{TopK}(\mathcal{A}_{k}\cap\mathcal{T})$ and top-K hidden state vision tokens $\bm{v}=\textit{TopK}(\mathcal{A}_{k}\cap\mathcal{V})$, where the top-K is ranked by the activation value $\bm{z}_{j}\bm{f}_{k}^{\top}$ of the token. We then compute the cosine similarity between the two lists of tokens, donating the cross-modal weight of feature $\bm{f}_{k}$ as
    $Cosine(\bm{t},\bm{v})=\frac{1}{k}\sum_{i=1}^{k}\frac{\bm{t}_{i}\cdot\bm{v}_{i}}{||\bm{t}_{i}||||\bm{v}_{i}||},$ (7)
@@ 3.1.1 Experiment Setup
For text-only and multimodal situations, we selected the Pile (Gao et al., 2020) and Obelics (Laurençon et al., 2023) datasets separately. Specifically, we sampled 100K data from each dataset as the train set and 10K data as the test set. The Pile is a diverse language modeling dataset for LLM pretraining, and Obelics is a massive interleaved image-text dataset for MLLM pretraining. These two datasets are widely recognized in various pretraining and interpretability works (Black et al., 2022; Biderman et al., 2023; Cunningham et al., 2023; Team, 2024).
@@
We selected two generic MLLMs, LLaVA-NeXT-7B (Liu et al., 2024) and Chameleon-7B (Team, 2024), as our target models. These two models represent two distinct architectures, and testing our method on them can prove that our method is applicable to different architectures.
@@ 3.1.2 Experiment Result
Table 1: The $\mathcal{L}_{0}$ metric of SAE and SAE-V models. $\mathcal{L}_{0}$ indicates the sparsity cost (average activated feature number). The results vary significantly across models due to their architectural differences. For Anole-7B and Chameleon-7B, SAE-V models maintain lower $\mathcal{L}_{0}$, suggesting more efficient feature utilization. However, LLaVA-NeXT-7B shows a contrary pattern with SAE-V model requiring higher feature activation than SAE. We propose that extra activated features of SAE-V model are introduced by extra vision tokens in multimodal data. Notably, Chameleon-7B with SAE model exhibits an unusually high sparsity cost, attributed to multiple unseen vision tokens in the inference stage.
@@
We compared the reconstruction performance of SAE-V model trained on LLaVA-NeXT-7B and SAE model to prove that SAE-V model trained on MLLMs can generalize to its base LLM. The findings (shown in Figure 4) indicate that across different settings, SAE-V model consistently achieves the best performance. Moreover, when trained on both MLLM and LLM, SAE model exhibits nearly identical reconstruction loss values, showing its robust transferability.
@@ 3.2 Apply SAE-V Model on Multimodal Data
The quantized results are presented in Figure 6, where we observe that, for all methods, preserving 75% or 50% of the patches achieves an accuracy close to that obtained using the full image. In the challenging scenario where only 25% of the patches are retained, $\mathcal{L}_{0}$ method, $\mathcal{L}_{1}$ method, and cosine similarity score method maintain accuracy levels close to 70%, and all methods significantly surpass the accuracy obtained by the random preservation method. These results demonstrate that SAE-V can accurately capture critical information in images and that the methods proposed in Section 2.2 effectively utilized SAE-V features during inference.
@@ 4.1 Experiment Setup
Consistent with Section 3.1.1, we selected LLaVA-NeXT-7B (Liu et al., 2024) and Chameleon-7B (Team, 2024) for our alignment experiment. Since the LLaVA-NeXT-7B model is rather powerful in multimodal capabilities, we selected the Align-Anything (Ji et al., 2024) text-image-to-text dataset for our experiment. Align-Anything is a 400K multimodal preference dataset containing fine-grained annotated multimodal input-output preference data, and we used the 40K subset of text-image input and text output in our experiment.
@@
To evaluate the efficiency of our methods, we applied Direct Preference Optimization (DPO) to the model using the filtered datasets.  [^5] We present detailed training parameters in Appendix B.2. We then evaluate the multimodal capabilities of the model using LLaVA-Bench (Liu et al., 2024) benchmarks.

### v2

@@ Abstract
With the integration of image modality, the semantic space of multimodal large language models (MLLMs) is more complex than text-only models, making their interpretability more challenging and their alignment less stable, particularly susceptible to low-quality data, which can lead to inconsistencies between modalities, hallucinations, and biased outputs. As a result, developing interpretability methods for MLLMs is crucial for improving alignment quality and efficiency. In text-only LLMs, Sparse Autoencoders (SAEs) have gained attention for their ability to interpret latent representations. However, extending SAEs to multimodal settings presents new challenges due to modality fusion and the difficulty of isolating cross-modal representations. To address these challenges, we introduce SAE-V, a mechanistic interpretability framework that extends the SAE paradigm to MLLMs. By identifying and analyzing interpretable features along with their corresponding data, SAE-V enables fine-grained interpretation of both model behavior and data quality, facilitating a deeper understanding of cross-modal interactions and alignment dynamics. Moreover, by utilizing cross-modal feature weighting, SAE-V provides an intrinsic data filtering mechanism to enhance model alignment without requiring additional models. Specifically, when applied to the alignment process of MLLMs, SAE-V-based data filtering methods could achieve more than 110% performance with less than 50% data. Our results highlight SAE-V’s ability to enhance interpretability and alignment in MLLMs, providing insights into their internal mechanisms. ***Our codebase and model are released at Github and Huggingface.
@@ 1 Introduction
With the development and success of large language models (LLMs) (Dubey et al., 2024; Achiam et al., 2023), researchers have begun to introduce visual understanding to these models, thereby extending their operational scope from language to a mix of vision and language, resulting in the creation of powerful multimodal large language models (MLLMs) (Alayrac et al., 2022; Liu et al., 2024a; Team et al., 2024; Team, 2024). To enhance the multimodal understanding capabilities of MLLMs, the research community has explored various architectures, including using individual image/text encoders to encode cross-modal information into a joint representation space (Zhang et al., 2023; Liu et al., 2024a; Zhu et al., 2024; Wu et al., 2024c) and leveraging image tokenizers to transform all inputs into a unified token sequence (Team, 2024; Xie et al., 2024; Wu et al., 2024a; Wang et al., 2024). Despite the difference in the architectures of these models, their essential goal is the same: Fuse the text and image representation space into a joint multimodal semantic space.
As MLLMs continue to scale up in both size and capability, their interpretability and controllability remain a significant challenge (Zhang & Zhu, 2018; Ben Melech Stan et al., 2024). Currently, mechanistic interpretability techniques such as circuit analysis (Olsson et al., 2022) and dictionary learning with sparse autoencoders (Huben et al., 2024) are the most widely recognized approaches to interpreting LLMs. However, their application to MLLMs, especially in the context of cross-modal integration, has been limited. There is a pressing need for specialized tools and frameworks that can unravel the intricate workings of MLLMs.
Moreover, current interpretability efforts are focused mainly on interpreting models, rather than applying this interpretability to real alignment situations, which also makes it difficult to evaluate these methods effectively. Top-down approaches, such as Representation Engineering (Zou et al., 2023) and activation steering (Turner et al., 2023; Rimsky et al., 2024), can directly evaluate the control effects of interpretability methods through control or unlearning techniques. However, for bottom-up methods like circuit analysis (Wang et al., 2023), sparse autoencoders (SAEs) (Huben et al., 2024), and cross-coders, effective evaluation methods beyond loss reduction are limited. Based on the previous discussion, can we propose a bottom-up multimodal interpretability approach that can directly enhance the alignment process?
@@ 2.1 Preliminary: Sparse Autoencoder Paradigm
We adopt SAE-V (denoted as $\mathcal{S}_{\theta}$) architecture from the methodology proposed in (Bricken et al., 2023), which comprises an encoder and a feature dictionary $\mathcal{F}_{\theta}\text{:}\ \{\bm{f}_{k}\}_{k=1}^{\bm{n}}$ as a decoder. Let the input be denoted as $H\in\mathbb{R}^{\bm{l}\times\bm{m}}$, where $\bm{l}$ donates the number of input tokens and $\bm{m}$ donates the shape of hidden state token, the hidden state of a specific layer of a MLLM $\mathcal{M}_{\theta}$. The SAE-V encoding operation $\mathcal{S}_{\theta}(\cdot)$ is defined as
$Z=\text{ReLU}(H\times W_{\text{enc}}+b_{enc}),$ (1)
@@
$\mathcal{L}_{R}={||H-Z\times(\bm{f}_{1},\bm{f}_{2},\dots,\bm{f}_{n})^{\top}||}_{2}^{2}.$ (2)
@@
where $\mathcal{L}_{1}=||Z||_{1}$ adds a sparsity constraint to the learned features and $\lambda$ is a hyperparameter controlling the level of sparsity. The training results could also quantized by incorporating an additional sparsity constraint via $\mathcal{L}_{0}=||Z||_{0}$, which counts the number of nonzero elements in the learned features $Z$.
@@
        $\mathcal{A}_{k}\leftarrow\mathcal{A}_{k}\cup\{\bm{h}_{j}\text{:}\ \bm{h}_{j}\in H_{i},\bm{z}_{j}=\bm{e}_{j}Z_{i},\bm{z}_{jk}>\delta\}$
@@
     $F_{i}\leftarrow\{\bm{f}_{k}\text{: }\exists\ z_{j}\in Z_{i},\ \mathrm{s.t.}\ z_{j}\ \text{activates }\bm{f}_{k}\}$
@@ 2.2 Interpreting Multimodal Data with SAE-V
It has been previously demonstrated (Gao et al., 2025; Huben et al., 2024) that SAE can be employed to interpret how LLMs encode semantic information from these models. This feature motivates us to apply SAE-V to assess the quality of the data and thus facilitate data filter for alignment.
We adopt a cosine similarity score ranking algorithm for data filtering (shown in Algorithm 1). Let the multimodal training dataset be donated as $\mathcal{D}=\{\bm{d}_{i}\}_{i=1}^{p}$, where $\bm{d}_{i}=(u_{\bm{1}},u_{\bm{2}},\dots,u_{\bm{l}})$, $u_{\bm{j}}\in\mathcal{T}\lor u_{\bm{j}}\in\mathcal{V},\bm{j}=\bm{1},\bm{2},\dots,\bm{l}$, is a sequence of tokens from text vocabulary $\mathcal{T}$ and tokens from vision vocabulary $\mathcal{V}$. We acquire feature activation token $\bm{z}_{j}$ by MLLM forward and Equation 1, i.e.
@@
We define a SAE-V feature $\bm{f}_{k}$ is activated on $\bm{z}_{j}\ \textit{if}\ \bm{z}_{jk}>\delta$, where $\delta$ is activation bound. Correspondingly, we state that $\bm{f}_{k}$ is activated on $\bm{d}_{i}\ \textit{if}\ \exists\ \bm{z}_{j}\in Z_{i},\ \mathrm{s.t.}\ z_{j}\ \text{activates}\ \bm{f}_{k}$.
@@
    SAE-V Feature Weighting For each feature $\bm{f}_{k}$, we identify its top-K hidden state text tokens $\bm{t}=\textit{TopK}(\mathcal{A}_{k}\cap\mathcal{T})$ and top-K hidden state vision tokens $\bm{v}=\textit{TopK}(\mathcal{A}_{k}\cap\mathcal{V})$, where the top-K is ranked by the activation value $\bm{z}_{jk}$ of the token. We then compute the cosine similarity between the two lists of tokens, donating the cross-modal weight of feature $\bm{f}_{k}$ as
    $Cosine(\bm{t},\bm{v})=\mathbb{E}_{i,j}\frac{\bm{t}_{i}\cdot\bm{v}_{j}}{||\bm{t}_{i}||||\bm{v}_{j}||},$ (7)
@@ 3 Interpretability Analysis with SAE-V
Figure 3: Examples of interpretable features discovered by SAE-V. We presented examples of interpretable SAE-V features on LLaVA-NeXT-7B that demonstrate cross-modal semantic consistency. (a) Feature #44031 exhibits strong activation for ‘Doberman dogs’ across both text and image modalities, showing SAE-V’s ability to identify specific concepts with concrete physical meanings. (b) Feature #11105 captures the abstract concept of ‘Symmetry’ across different modalities, activating for various symmetry patterns including left-right, top-bottom, and central symmetry, with activation regions precisely aligned with symmetrical elements in images. These examples illustrate that SAE-V can discover features representing both concrete entities and abstract concepts that maintain consistent semantic meaning across modalities, going beyond what traditional probing methods on raw activations can achieve.
@@ 3.1.1 Experiment Setup
For text-only and multimodal situations, we selected the Pile (Gao et al., 2020) and Obelics (Laurençon et al., 2023) datasets separately. Specifically, we sampled 100K data from each dataset as the train set and 10K data as the test set. The Pile is a diverse language modeling dataset for LLM pretraining, and Obelics is a massive interleaved image-text dataset for MLLM pretraining. These two datasets are widely recognized in various pretraining and interpretability works (Black et al., 2022; Biderman et al., 2023; Huben et al., 2024; Team, 2024).
@@
We selected two generic MLLMs, LLaVA-NeXT models (including LLaVA-NeXT-Mistral-7B, LLaVA-NeXT-Vicuna-7B and LLaVA-NeXT-Vicuna-13B) (Liu et al., 2024a) and Chameleon-7B (Team, 2024), as our target models. These models represent two distinct architectures, and testing our method on them can demonstrate that our method is applicable to different architectures.
@@ 3.1.2 Experiment Result
LLaVA-NeXT-Mistral-7B SAE 94.5
SAE-V 192.5
LLaVA-NeXT-Vicuna-7B SAE 3162.96
SAE-V 585.64
LLaVA-NeXT-Vicuna-13B SAE 128.56
SAE-V 193.63
Table 1: The $\mathcal{L}_{0}$ metric of SAE and SAE-V models. $\mathcal{L}_{0}$ indicates the sparsity cost (average activated feature number). The results vary significantly across models due to their architectural differences. For Anole-7B and Chameleon-7B, SAE-V models maintain lower $\mathcal{L}_{0}$, suggesting more efficient feature utilization. However, LLaVA-NeXT models shows a contrary pattern with SAE-V requiring higher feature activation than SAE. We propose that extra activated features of SAE-V model are introduced by extra vision tokens in multimodal data. Notably, Chameleon-7B and LLaVA-NeXT-Vicuna-7B with SAE model exhibits an unusually high sparsity cost, attributed to multiple unseen vision tokens in the inference stage.
@@
The transferability of SAEs between foundation models and instruction-tuned models has been extensively investigated in text-only contexts (Kissane et al., 2024; Kutsyk et al., 2024; Gallifant et al., 2025), as it demonstrates whether SAEs can capture universal semantic features within LLMs. Similarly, the transferability from MLLMs to corresponding LLMs serves as a critical metric for the quality of features learned by SAE-V. We compared the reconstruction performance of SAE-V model trained on LLaVA-NeXT-7B and SAE model to prove that SAE-V model trained on MLLMs can generalize to its base LLM. The findings (shown in Figure 5) indicate that across different settings, SAE-V model consistently achieves the best performance. Moreover, when trained on both MLLM and LLM, SAE model exhibits nearly identical reconstruction loss values, showing its robust transferability.
@@ 3.2 Apply SAE-V Model on Multimodal Data
The quantized results are presented in Figure 7, where we observe that, for all methods, preserving 75% or 50% of the patches achieves an accuracy close to that obtained using the full image. In the challenging scenario where only 25% of the patches are retained, $\mathcal{L}_{0}$ method, $\mathcal{L}_{1}$ method, and cosine similarity score method maintain accuracy levels close to 70%, and all methods significantly surpass the accuracy obtained by the random preservation method. These results demonstrate that SAE-V can accurately capture critical information in images and that the methods proposed in Section 2.2 effectively utilize SAE-V features during inference.  [^5] More experiments and their quantative analysis are presented in Appendix  C.3
@@ 4.1 Experiment Setup
(a) Performance on Align-Anything dataset
(b) Performance on MMInstruct dataset
(c) Performance on RLAIF-V dataset
Figure 8: Ablation study of SAE-V-based data filter method on Chameleon-7B model and different datasets. We replicate SAE-V-based data filter on Chameleon-7B model and three distinct datasets. The results show that SAE-V-based data filter method is still effective on different datasets and different model architectures and image encoding methods other than CLIP. (a) On Align-Anything dataset, both cosine similarity and co-occurrence filters perform better than random filters at almost every data percentage, and the cosine similarity filter achieved a score of 52.1 (120% of the full dataset’s performance) with 70% data. (b) On MMInstruct, the cosine similarity filter outperforms the random filter baseline at every data proportions. Specifically, the cosine similarity filter achieved a score of 61.2 (117% of the full dataset’s performance) with 60% data. (c) On RLAIF-V dataset, the cosine similarity filter achieves the highest score of 54.2 (125% of the full dataset’s performance) with only 10% data, demonstrating its supreme efficiency.
@@
Consistent with Section 3.1.1, we selected LLaVA-NeXT-7B (Liu et al., 2024a) and Chameleon-7B (Team, 2024) for our alignment experiment. Since the LLaVA-NeXT-7B model is rather powerful in multimodal capabilities, we selected the Align-Anything (Ji et al., 2024) text-image-to-text dataset for our experiment. Align-Anything is a 400K multimodal preference dataset containing fine-grained annotated multimodal input-output preference data, and we used the 40K subset of text-image input and text output in our experiment.
@@
To evaluate the efficiency of our methods, we applied Direct Preference Optimization (DPO) to the model using the filtered datasets.  [^7] We present detailed training parameters in Appendix B.2. We then evaluate the multimodal capabilities of the model using LLaVA-Bench (Liu et al., 2024a) benchmarks.
Figure 9: The performance of SAE-V-based data filter method We evaluated SAE-V-based data filter method on LLaVA-NeXT-7B model, Align-Anything dataset, and LLaVA-Bench benchmark. The results show that all SAE-V-based methods significantly outperform the random selection baseline, while the cosine similarity filter achieved 108% of the full dataset’s performance with only 20% of the data, and the co-occurrence filter peaked at 50% of the data, reaching a score of 108.17. As a more straightforward utilization of SAE-V model, the L0 filter also generally outperforms the random selection baseline.